# Classifying Infinity Topoi via weighted limits

### v1

@@ Introduction
The classifying topos $\mathsf{Set}[\mathbb{T}]$ of a geometric theory $\mathbb{T}$ provides access to the so called generic model of a theory, which controls the generic behavior of the whole theory.
Besides many applications to logic (see part D of [Joh02]), some early applications to algebraic geometric were observed by Wraith [Wra06], and later displayed by Blechschmidt [Ble21, Ble18], whose work has influenced and inspired many in the last years [Ble20, CCH24, CCRW24] in the general topic of synthetic algebraic geometry. Many other applications of this technology to other fields of mathematics have been provided by Caramello [Car23, Car16].
@@
This paper deals with the general problem of building classifying $\infty$-topoi. There are several motivations to approach this problem, some very concrete and some very conceptual. Many of them can be understood as a homotopical-replay of the classical motivations, investing several topics of pure mathematics. Others protrude to the question of what homotopical geometric logic is supposed to be, and thus pertain more to logic. The question concerning the existence of some specific classifying $\infty$-topoi has appeared a number of times in the community of homotopical algebra [hh, hr].
In the last years there have been some attempts to define a satisfactory notion of $\infty$-geometric theory. Prominently, Stenzel [Ste22, Ste23] has provided the best results so far, introducing notions of regular, coherent and geometric $\infty$-theories. Unfortunately, these notions are not supported by a proof calculus which is as robust as that of geometric logic, and require a much more sophisticated treatment than the usual $1$-dimensional counterpart. The main obstruction to such development at the moment seems to be conceptual. An $\infty$-topos has an object classifier and the complexity of its localizations cannot be reduced to its classifier of subobjects. This is mirrored by the complexity of the theory of sites, which requires the more flexible notion of lex-modulator [ABFJ22, ABFJ24] to be developed, while the more familiar notion of $\infty$-site by Lurie [Lur09a] only recovers hypercomplete topoi. Currently it is unclear (and to some extent ultimately open to debate) whether a compact notion of $\infty$-geometric theory can be defined without the complexity of Stenzel, and in the quest for defining classifying topoi, the question classifying topoi of what? remains somewhat unanswered.
@@
In this paper we suspend our judgment concerning the notion of geometric $\infty$-theory and take a more concrete approach to build classifying topoi. In [Joh02, part B], Johnstone provides an alternative approach to the construction of classifying topoi which precedes the notion of geometric theory, and for every intended application is indistinguishable from its more logical counterpart. That is the construction of classifying topoi via weighted limits. From a technical point of view, there are two main results in the paper. The first one concerns the $(\infty,2)$-category of topoi.
Theorem 0.1 (Theorem 5.1).
@@
The second one concerns the existence of classifying topoi. Indeed in the last section we deploy this technology to construct classifying topoi. For us, a theory will just be a construction – or to be precise a prestack –,
@@
which should be identified with the association mapping a topos $\mathcal{E}$ to the category of models of the intended theory $\mathbb{T}$ in $\mathcal{E}$. Of course, since no logic has been defined, this is just an informal suggestion, and this notion of theory is indeed much more general than an appropriate notion of geometric theory would be. Being more general though, this notion captures a very big number of interesting constructions. For a theory to have a classifying topos means that the prestack is represented by a topos,
$\mathbb{T}\simeq TOP_{\mathcal{B}}(-,\mathcal{B}).$
Example 0.2 (Models fo a Lawvere theory).
Let $\mathsf{L}$ be a Lawvere theory in the sense of [Ber20], then we have a canonically associated theory mapping a topos $\mathcal{E}$ to its $\infty$-category of $\mathsf{L}$-models in $\mathcal{E}$, $\mathbb{T}^{\mathsf{L}}:TOP_{\mathcal{S}}^{op}\to CAT_{\infty}.$
Example 0.3 (Example 7.9).
For $\mathcal{E}$ a topos we can construct the category of spectra objects in $\mathcal{E}$ and consider the associated prestack, $\mathbb{S}:TOP_{\mathcal{S}}^{op}\to CAT_{\infty}.$
We show that these two prestacks both have a classifying topos and we provide a general recipe to construct classifying topoi of geometric sketches. The notion of sketch was recently imported to $\infty$-categories in [CGMC25] and has proven to be extremely expressive, overarching essentially every imaginable construction. In our paper we slightly restric the focus to geometric sketches (which can be taken as a practical notion of geometric theory) and prove:
Theorem 0.4 (Theorem 7.12).
Let $\Sigma$ be a geometric sketch, then the associated prestack of models has a classifying topos.
@@
As it is evident from the introduction, most of our effort is devoted to proving Theorem 5.1. To do so we essentially reproduce the original proof-strategy by Johnstone. In Section 1 we establish the notation on $\infty$-categories and $(\infty,2)$-categories). In Section 2 we show that an $(\infty,2)$-category has weighted limits if and only if it has conical limits and powers with the arrow (Corollary 2.7). It was already observed by Lurie that the $(\infty,1)$-category of topoi has all limits, and hence we are deduced to prove that it we can power along the arrow. In Section 3 we briefly review the rudiments of internal higher category theory, as developed in [Mar22], [MW24] and [MW25]. In Section 4 we prove that indeed we can power via the arrow, and finally in Section 5 we establish the appropriate bicategory of topoi and prove Theorem 5.1. In Section 6 we construct the object classifier, which is the blueprint for the whole theory of classifying topoi. Finally in Section 7 we discuss the notion of classifying topos of a geometric sketch and provide many examples.
@@ Acknowledgments
The first named author received funding from Knut and Alice Wallenberg Foundation, grant no. 2020.0199. The second named author was author was supported by the Starting Grant 101077154 “Definable Algebraic Topology” from the European Research Council awarded to Martino Lupini.
@@ 1.1. Background on $\infty$-categories
In working with $\infty$-categories, we will often work in the context of Jacob Lurie’s book [Lur09a] -see there for more details. An $\infty$-category is by definition a quasi-category, i.e. a simplicial set satisfying the appropriate lifting property. We often will write objects (or 0-simplices) in an $\infty$-category by lower case letters, such as $x,y$. We call the 1-simplices of an $\infty$-category edges or 1-morphisms. An edge is said to be an equivalence if and only if it represents an equivalence in the homotopy category of an $\infty$-category (see [Lur09a, Section 1.2.3] for the definition of the homotopy category). An $\infty$-category is called an $\infty$-groupoid, iff it each edge is invertible. In this case, the $\infty$-groupoids are precisely the Kan complexes by a theorem of Joyal ([Joy02]).
@@
Given two objects $x,y$ in an $\infty$-category $C$, we will write $Hom_{C}(x,y)$ for the space of maps between $x$ and $y$. We will be working in a relatively model-independent manner, so it does not matter which of the (equivalent) models of mapping spaces from [Lur09a, Section 1.2.2] we use. An equivalence of $\infty$-categories is just an equivalence in Joyal’s model structure for $\infty$-categories. That is, it induces an equivalence of homotopy categories, as well as induces weak equivalences of mapping spaces.
@@ 1.2. Background on $\infty$-bicategories
$Set_{\Delta}^{sc}\rightarrow Set_{\Delta}$
@@
Given a scaled simplicial set $X$, we define its core to be the simplicial set $X^{th}$ spanned by those n-simplices of $X$ whose 2-dimensional faces are thin triangles (note that here $th$ stands for thin). The assignment $X\mapsto X^{th}$ is right adjoint to the forgetful functor $(-)_{\sharp}$
@@
We say that a scaled simplicial set $C$ is an $\infty$-bicategory if the map $C\rightarrow*$ has the right lifting property with respect to the class of maps $\mathbf{S}$ from [GHL21, Definition 1.2.6]
@@
As noted in [Lur09b], the cartesian product is a left Quillen bifunctor. In particular, for any two scaled simplicial sets $X,Y$, we have a mapping object $Fun(X,Y)$, which satisfies,
$Hom_{Set_{\Delta}^{sc}}(Z,Fun(X,Y))\cong Hom_{Set_{\Delta}^{sc}}(Z\times X,Y)$
@@
$\mathfrak{C}^{sc}:sSet^{sc}_{\Delta}\leftrightarrows sSet^{+}_{\Delta}:N^{sc}$
where the right adjoint is called the scaled coherent nerve. Applying the scaled coherent to the full subcategory $sSet^{+}$ spanned by $\infty$-categories, we obtain the $\infty$-bicategory of $\infty$-categories, which we denote by $CAT_{\infty}$. Note that $CAT_{\infty}^{th}=Cat_{\infty}$.
@@
Suppose that $C$ is an $\infty$-bicategory. Then for each $x,y\in C_{0}$ will write $Hom_{C}(x,y)$ for the simplicial set whose n-simplices are maps $s:\Delta^{n}\times\Delta^{1}\rightarrow C$ such that $s|_{\Delta^{n}\times\{0\}}=x,s|_{\Delta^{n}\times\{1\}}=y$ and $s|_{\Delta^{\{(i,0),(i,1),(j,1)\}}}$ is thin.
@@
Suppose that $\mathcal{C}$ is an $\infty$-bicategory. Then
@@
    $Hom_{C}(x,y)$ is an $\infty$-category for each $x,y\in C$
@@
    We have an identification $\mathbf{i}Hom_{C}(x,y)=Hom_{C^{th}}(x,y)$
@@
For the second statement, we first want to show that there is a natural inclusion $\mathbf{i}Hom_{C}(x,y)\subseteq Hom_{C^{th}}(x,y)$, where $Hom_{\mathcal{C}^{th}}$ is the model of mapping space in an $\infty$-category described in [Lur09a, pg. 28]. To do this, it suffices to show that each map $s:\Delta^{n}\times\Delta^{1}\rightarrow C$ representing an n-simplex of $\mathbf{i}Hom_{C}(x,y)$ factors through $C^{th}$. In other words, we want to show that each 2-simplex of $\Delta^{n}\times\Delta^{1}$ maps under $s$ to a thin 2-simplex. Since each degenerate 2-simplex is automatically, thin it suffices to show that any map $e:=\Delta^{1}\times\Delta^{1}\subseteq\Delta^{n}\times\Delta^{1}\xrightarrow{F}C$ corresponding to an edge $\mathbf{i}Hom_{C}(x,y)$, maps triangles to thin triangles. By definition, $e|_{\Delta^{1}\times 0},e|_{\Delta^{1}\times\{1\}}$ are both equivalences. Thus, by the construction of mapping spaces from [Lur09a, Section 1.2.2] , $e$ maps each edge of $\Delta^{1}\times\Delta^{1}$ to equivalences. Since triangles spanned by equivalences are thin in an $\infty$-bicategory, we have that $e$ factors through $C^{th}$ as required.
@@
(see [GHL21, Definition 4.2.1]) Suppose that $C$ is an $(\infty,2)$-category and $x\in\mathcal{C}$. Then the slice $C^{/x}$ is an $(\infty,2)$-category such that the n-simplices of $(C^{/x})^{th}$ can be identified with maps $\phi:\Delta^{1}\times\Delta^{n}\rightarrow C^{th}$ with $\phi|_{\Delta^{\{1\}\times\Delta^{n}}}=x$ and the thin simplices can be identified with maps
$\Delta^{1}_{\flat}\otimes\Delta^{2}_{\sharp}\rightarrow\mathcal{C}$
where $\otimes$ is the Gray tensor product of [GHL21, Definition 4.1.1] Thus, by the definition of the gray tensor product, there is a natural inclusion $C^{/x}\rightarrow Fun(\Delta^{1}_{\flat},C)$.
@@ 2. $(\infty,2)$-Categorical Limits from Conical Limits and Arrows
Let $C$ be an $\infty$-bicategory and $f:J\rightarrow C$ and $w:J\rightarrow CAT_{\infty}$ be two functors of $\infty$-bicategories. Then the limit $l\in C$ of $f$ weighted by $w$ is characterized by the natural equivalence of $\infty$-categories.
(1) $Hom_{C}(x,l)\cong Nat_{J}(w,Hom_{C}(x,f(-)))$
where the right-hand side denotes the mapping category in the $(\infty,2)$-bicategory $Fun(J,CAT_{\infty})$ between $w$ and $Hom_{C}(x,f(-)))$.
@@
Suppose that $X$ is a cocomplete $(\infty,1)$-category. Then $X$ is cotensored over $Cat_{\infty}$ if it admits a cotensoring with $\Delta^{1}$.
@@
Consider the full subcategory $\{\Delta^{1}\}\subseteq Cat_{\infty}$ on the interval $\Delta^{1}$. We will define the cotensoring to be the left Kan extension
${\{\Delta_{1}\}\times X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta^{1}\otimes(-)}$${X}$${Cat_{\infty}\times X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-)\otimes(-)}$
By [Lur09a, Lemma 4.3.2.13] and the cocompleteness of $X$, the left Kan extension exists.
By [Lur09a, Proposition 4.2.3.8], we can regard as $F$ as the left Kan extension of $(-)\otimes(-)|_{\Delta\times X}$. Note that by [Hau22, Lemma 2.5], $\Delta$ is a dense subcategory of $Cat_{\infty}$ in the sense that for each $\in Cat_{\infty}, xx$ is the colimit of
@@
$Hom_{X}(c\otimes\Delta^{n},d)\cong Hom_{Cat_{\infty}}(\Delta^{n},Hom_{X}(c,d))$
@@
However, appealing to [Lur09a, Proposition 4.2.3.8] again, we note that $(-)\otimes(-)|_{\Delta\times X}$ is the left Kan extension of $\Delta^{1}\times(-)$ along $\{\Delta^{1}\}\subseteq\Delta$. Thus, reasoning as before, it suffices to show that $\Delta^{1}$ is dense in $\Delta$. But the fact that $\Delta^{n}$ is the colimit of
@@
For brevity, let $M_{x}=Hom_{C}(x,l),N_{x}=Nat_{J}(w,Hom_{C}(x,F(-)))$. Let $a,b\in M_{x}$. Consider the commutative daigram
@@
Suppose that $C$ is an $\infty$-bicategory that has all limits indexed by $\infty$-categories, and a functor $(-)^{\Delta^{1}}$. Then $X$ is cotensored over $CAT_{\infty}$.
@@
$x^{C}=\underset{(\{\Delta^{1}\}_{/C})\xrightarrow{p}CAT_{\infty}}{colim}x^{\Delta^{1}}$
$(-)_{/C}$ is the slice construction from [GHL21, Section 4.1] and $p$ is the projection. We need to verify that this tensor product satisfies the appropriate universal property. To prove the property in general, we can assume that $C=CAT_{\infty}$, by the representable nature of limits. In this case, since $CAT_{\infty}$ has cotensors with $\Delta^{1}$, we can use Lemma 2.4 to show that the underlying functor of $\infty$-categories satisfies the appropriate universal property. However, the underlying $\infty$-category of $C_{/X}$ is $(C^{th})_{/X}$ by the construction of [GHL21, Section 4.1]. Thus, the construction reduces to the construction of Lemma 2.3 on underlying $\infty$-categories, hence the result.
@@
Suppose that $C$ is an $\infty$-bicategory. Then if $C$ has a powering $(-)^{\Delta^{1}}$ and $C^{th}$ has all limits, then $C$ has all limits
@@
By the representable nature of limits, we are reduced to proving the statement for $CAT_{\infty}$. In this case, by Lemma 2.4, we want to show that weighted limits in the underlying $\infty$-category can be computed as ordinary limits. Since $C^{th}$ is cotensored over $Cat_{\infty}$ this follows from the formulas for weighted colimits from [GHN17] in terms of the twisted arrow category.
@@
Suppose that $C$ is an $\infty$-category. Suppose that $C^{th}$ has all finite limits and a powering $(-)^{\Delta^{1}}$. Then $C$ has all finite weighted limits.
@@
Since $\Delta_{/J}$ is finite when $J$ is, finite weighted limits exist in the presence of ordinary finite limits and cotensors with finite $\infty$-categories.

### v2

@@ Introduction
The classifying topos $\mathsf{Set}[\mathbb{T}]$ of a geometric theory $\mathbb{T}$ provides access to the so-called generic model of a theory, which controls the generic behavior of the whole theory.
Besides many applications to logic (see part D of [Joh02]), some early applications to algebraic geometry were observed by Wraith [Wra06], and later displayed by Blechschmidt [Ble21, Ble18], whose work has influenced and inspired many other researchers in the last years [Ble20, CCH24, CCRW24] in the general topic of synthetic algebraic geometry. Many other applications of this technology to other fields of mathematics have been provided by Caramello [Car23, Car16].
@@
This paper deals with the general problem of building classifying $\infty$-topoi. There are several motivations to approach this problem, some very concrete and some very conceptual. Many of them can be understood as a homotopical replay of the classical motivations, investing several topics of pure mathematics. Others protrude to the question of what homotopical geometric logic is supposed to be, and thus pertain more to logic. The question concerning the existence of some specific classifying $\infty$-topoi has appeared a number of times in the community of homotopical algebra [hh, hr] and it was very informally addressed in the last section of [ABFJ22].
In the last years there have been some attempts to define a satisfactory notion of $\infty$-geometric theory. Prominently, Stenzel [Ste22, Ste23] has provided the best results so far, introducing notions of regular, coherent and geometric $\infty$-theories. Unfortunately, these notions are not supported by a proof calculus which is as robust as that of geometric logic, and require a much more sophisticated treatment than the usual $1$-dimensional counterpart. The main obstruction to such development at the moment seems to be conceptual. An $\infty$-topos has an object classifier and the complexity of its localizations cannot be reduced to its classifier of subobjects. This is mirrored by the complexity of the theory of sites, which requires the more flexible notion [ABFJ22, ABFJ24] to be developed, while the more familiar notion of $\infty$-site by Lurie [Lur09a] only recovers hypercomplete topoi. Currently it is unclear (and to some extent ultimately open to debate) whether a compact notion of $\infty$-geometric theory can be defined without the complexity of Stenzel, and in the quest for defining classifying topoi, the question classifying topoi of what? remains somewhat unanswered.
@@
In this paper, we suspend our judgment concerning the notion of geometric $\infty$-theory and take a more concrete approach to build classifying topoi. In [Joh02, part B], Johnstone provides an alternative approach to the construction of classifying topoi which precedes the notion of geometric theory, and for every intended application is indistinguishable from its more logical counterpart. That is, the construction of classifying topoi via weighted limits. From a technical point of view, there are two main results in the paper. The first one concerns the $(\infty,2)$-category of topoi.
Theorem (Theorem 5.1).
@@
The second main result is about the existence of classifying topoi. Indeed, in the last section we deploy this technology to construct classifying topoi. For us, a theory will just be a construction, or more precisely, a prestack
@@
which should be identified with the association mapping a topos $\mathcal{E}$ to the category of models of the intended theory $\mathbb{T}$ in $\mathcal{E}$. Of course, since no logic has been defined, this is just an informal suggestion, and this notion of theory is indeed much more general than an appropriate notion of geometric theory would be. Being more general, this notion captures a very large number of interesting constructions. For a theory to have a classifying topos means that the prestack is represented by a topos,
$\mathbb{T}\simeq TOP_{\mathcal{B}}(-,\mathcal{B}[\mathbb{T}]).$
Example (Models of a Lawvere theory).
Let $\mathsf{L}$ be a Lawvere theory in the sense of [Ber20]. Then we have a canonically associated theory mapping a topos $\mathcal{E}$ to its $\infty$-category of $\mathsf{L}$-models in $\mathcal{E}$, $\mathbb{T}^{\mathsf{L}}:TOP_{\mathcal{S}}^{op}\to CAT_{\infty}.$
Example (Example 7.9).
For $\mathcal{E}$ a topos we can construct the category of spectrum objects in $\mathcal{E}$ and consider the associated prestack, $\mathbb{S}:TOP_{\mathcal{S}}^{op}\to CAT_{\infty}.$
We show that these two prestacks both have a classifying topos and we provide a general recipe to construct classifying topoi of geometric sketches. The notion of sketch was recently imported to $\infty$-categories in [CGMC25] and has proven to be extremely expressive, overarching essentially every imaginable construction. In our paper we slightly restrict the focus to geometric sketches (which can be taken as a practical notion of geometric theory) and prove the following.
Theorem (Theorem 7.13 and Theorem 7.17).
Let $\mathbb{T}$ be a theory, then the following are equivalent:
  • $\mathbb{T}$ admits a classifying topos;
  • $\mathbb{T}$ is the prestack of models of a (small) geometric sketch.
@@
As it is evident from the introduction, most of our effort is devoted to proving Theorem 5.1. To do so we essentially reproduce the original strategy for the proof from Johnstone. In Section 1 we establish the notation on $\infty$-categories and $(\infty,2)$-categories). In Section 2 we show that an $(\infty,2)$-category has weighted limits if and only if it has conical limits and powers with the arrow (Corollary 2.7). It was already observed by Lurie that the $(\infty,1)$-category of topoi has all limits, and hence we are deduced to prove that we can power along the arrow. In Section 3, we briefly review the rudiments of internal higher category theory, as developed in [Mar22], [MW24] and [MW25]. In Section 4, we prove that indeed we can power via the arrow, and finally in Section 5 we establish the appropriate $\infty$-bicategory of topoi and prove Theorem 5.1. In Section 6, we construct the object classifier, which is the blueprint for the whole theory of classifying topoi. Finally, in Section 7, we discuss the notion of classifying topos of a geometric sketch and provide many examples.
@@ Acknowledgments
The first named author received funding from Knut and Alice Wallenberg Foundation, grant no. 2020.0199. The second named author was supported by the Starting Grant 101077154 “Definable Algebraic Topology” from the European Research Council awarded to Martino Lupini.
@@ 1.1. Background on $\infty$-categories
In working with $\infty$-categories, we will often work in the context of Jacob Lurie’s book [Lur09a] -see there for more details. An $\infty$-category is by definition a quasi-category, i.e. a simplicial set satisfying the appropriate lifting property. We often will write objects (or 0-simplices) in an $\infty$-category by lower case letters, such as $x,y$. We call the 1-simplices of an $\infty$-category edges or 1-morphisms. An edge is said to be an equivalence if and only if it represents an equivalence in the homotopy category of an $\infty$-category (see [Lur09a, Section 1.2.3] for the definition of the homotopy category). An $\infty$-category is called an $\infty$-groupoid, iff each of its edges are invertible. In this case, the $\infty$-groupoids are precisely the Kan complexes by a theorem of Joyal ([Joy02]).
@@
Given an object $x$ in an $\infty$-category $C$, we will write $C_{/x}$ and $C^{/x}$ for the slice constructions of [Lur09a, Proposition 1.2.9.2 and Definition 4.2.1.1], respectively. We will write $C_{x/}$ and $C^{x/}$ for their dual coslice constructions. Given two objects $x,y$ in an $\infty$-category $C$, we will write $Hom_{C}(x,y)$ for the space of maps between $x$ and $y$. We will be working in a relatively model-independent manner, so it does not matter which of the (equivalent) models of mapping spaces from [Lur09a, Section 1.2.2] we use. An equivalence of $\infty$-categories is just an equivalence in Joyal’s model structure for $\infty$-categories. That is, it induces an equivalence of homotopy categories, as well as induces weak equivalences of mapping spaces.
@@ 1.2. Background on $\infty$-bicategories
$Set_{\Delta}^{sc}\rightarrow Set_{\Delta}.$
@@
Given a scaled simplicial set $X$, we define its core to be the simplicial set $X^{th}$ spanned by those n-simplices of $X$ whose 2-dimensional faces are thin triangles (note that here $th$ stands for thin). The assignment $X\mapsto X^{th}$ is right adjoint to the forgetful functor $(-)_{\sharp}$.
@@
We say that a scaled simplicial set $C$ is an $\infty$-bicategory if the map $C\rightarrow*$ has the right lifting property with respect to the class of maps $\mathbf{S}$ from [GHL21, Definition 1.2.6].
@@
As noted in [Lur09b], the cartesian product is a left Quillen bifunctor. In particular, for any two scaled simplicial sets $X,Y$, we have a mapping object $Fun(X,Y)$, which satisfies
$Hom_{Set_{\Delta}^{sc}}(Z,Fun(X,Y))\cong Hom_{Set_{\Delta}^{sc}}(Z\times X,Y).$
@@
$\mathfrak{C}^{sc}:Set^{sc}_{\Delta}\leftrightarrows Set^{+}_{\Delta}:N^{sc}$
where the right adjoint is called the scaled coherent nerve. Applying the scaled coherent nerve to the full subcategory $Set^{+}_{\Delta}$ spanned by $\infty$-categories, we obtain the $\infty$-bicategory of $\infty$-categories, which we denote by $CAT_{\infty}$. Note that $CAT_{\infty}^{th}=Cat_{\infty}$.
@@
Suppose that $C$ is an $\infty$-bicategory. Then for each $x,y\in C_{0}$, we will write $Hom_{C}(x,y)$ for the simplicial set whose $n$-simplices are maps $s:\Delta^{n}\times\Delta^{1}\rightarrow C$ such that $s|_{\Delta^{n}\times\{0\}}=x,s|_{\Delta^{n}\times\{1\}}=y$ and $s|_{\Delta^{\{(i,0),(i,1),(j,1)\}}}$ is thin.
@@
Suppose that $\mathcal{C}$ is an $\infty$-bicategory. Then we have the following:
@@
    $Hom_{C}(x,y)$ is an $\infty$-category for each $x,y\in C$.
@@
    We have an identification $\mathbf{i}Hom_{C}(x,y)=Hom_{C^{th}}(x,y)$.
@@
For the second statement, we first want to show that there is a natural inclusion $\mathbf{i}Hom_{C}(x,y)\subseteq Hom_{C^{th}}(x,y)$, where $Hom_{\mathcal{C}^{th}}$ is the model of mapping space in an $\infty$-category described in [Lur09a, pg. 28]. To do this, it suffices to show that each map $s:\Delta^{n}\times\Delta^{1}\rightarrow C$ representing an $n$-simplex of $\mathbf{i}Hom_{C}(x,y)$ factors through $C^{th}$. In other words, we want to show that each 2-simplex of $\Delta^{n}\times\Delta^{1}$ maps under $s$ to a thin 2-simplex. Since each degenerate 2-simplex is automatically thin, it suffices to show that any map $e:=\Delta^{1}\times\Delta^{1}\subseteq\Delta^{n}\times\Delta^{1}\xrightarrow{F}C$ corresponding to an edge $\mathbf{i}Hom_{C}(x,y)$, maps triangles to thin triangles. By definition, $e|_{\Delta^{1}\times\{0\}},e|_{\Delta^{1}\times\{1\}}$ are both equivalences. Thus, by the construction of mapping spaces from [Lur09a, Section 1.2.2] , $e$ maps each edge of $\Delta^{1}\times\Delta^{1}$ to equivalences. Since triangles spanned by equivalences are thin in an $\infty$-bicategory, we have that $e$ factors through $C^{th}$ as required.
@@
(see [GHL21, Definition 4.2.1]) Suppose that $C$ is an $\infty$-bicategory and $x\in\mathcal{C}$. Then the slice $C^{/x}$ is an $\infty$-bicategory such that the n-simplices of $(C^{/x})^{th}$ can be identified with maps $\phi:\Delta^{1}\times\Delta^{n}\rightarrow C^{th}$ with $\phi|_{\{1\}\times\Delta^{n}}=x$ and the thin simplices can be identified with maps
$\Delta^{1}_{\flat}\otimes\Delta^{2}_{\sharp}\rightarrow C$
where $\otimes$ is the Gray tensor product of [GHL21, Definition 4.1.1]. Thus, by the definition of the Gray tensor product, there is a natural inclusion $C^{/x}\rightarrow Fun(\Delta^{1}_{\flat},C)$.
@@ 2. $(\infty,2)$-Categorical Limits from Conical Limits and Arrows
Let $C$ be an $\infty$-bicategory and $f:J\rightarrow C$ and $w:J\rightarrow CAT_{\infty}$ be two functors of $\infty$-bicategories. Then the limit $l\in C$ of $f$ weighted by $w$ is characterized by the natural equivalence of $\infty$-categories:
(1) $Hom_{C}(x,l)\cong Nat_{J}(w,Hom_{C}(x,f(-))),$
where the right-hand side denotes the mapping category in the $\infty$-bicategory $Fun(J,CAT_{\infty})$ between $w$ and $Hom_{C}(x,f(-)))$.
@@
Suppose that $X$ is a complete $\infty$-category. Then $X$ is cotensored over $Cat_{\infty}$ if it admits a cotensoring with $\Delta^{1}$.
@@
Consider the full subcategory $\{\Delta^{1}\}\subseteq Cat_{\infty}$ on the interval $\Delta^{1}$. We will define the cotensoring to be the right Kan extension
${(\{\Delta_{1}\})^{op}\times X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-)^{\Delta^{1}}}$${X}$${Cat_{\infty}^{op}\times X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-)^{(-)}}$
By [Lur09a, Lemma 4.3.2.13] and the completeness of $X$, the right Kan extension exists.
By [Lur09a, Proposition 4.2.3.8], we can regard as $F$ as the right Kan extension of $(-)^{(-)}|_{\Delta\times X}$. Note that by [Hau22, Lemma 2.5], $\Delta$ is a dense subcategory of $Cat_{\infty}$ in the sense that for each $\in Cat_{\infty}, xx$ is the colimit of
@@
$Hom_{X}(c,d^{\Delta^{n}})\cong Hom_{Cat_{\infty}}(\Delta^{n},Hom_{X}(c,d))$
@@
However, appealing to [Lur09a, Proposition 4.2.3.8] again, we note that $(-)^{(-)}|_{\Delta\times X}$ is the right Kan extension of $(-)^{\Delta^{1}}$ along $\{\Delta^{1}\}\subseteq\Delta$. Thus, reasoning as before, it suffices to show that $\Delta^{1}$ is dense in $\Delta$. But the fact that $\Delta^{n}$ is the colimit of
@@
For brevity, let $M_{x}=Hom_{C}(x,l),N_{x}=Nat_{J}(w,Hom_{C}(x,F(-)))$. Let $a,b\in M_{x}$. Consider the commutative diagram
@@
Suppose that $C$ is an $\infty$-bicategory that has all limits indexed by $\infty$-categories, and a functor $(-)^{\Delta^{1}}:C\rightarrow C$. Then $X$ is cotensored over $CAT_{\infty}$.
@@
$x^{C}=\underset{(\{\Delta^{1}\}^{/C})\xrightarrow{p}CAT_{\infty}}{lim}x^{\Delta^{1}}$
$(-)^{/C}$ is the slice construction from 1.19 $p$ is the projection. We need to verify that this cotensor product satisfies the appropriate universal property. By the representable nature of limits and the fact that $CAT_{\infty}$ has tensors with $\Delta^{1}$, Lemma 2.4 implies that it suffices to show that the underlying functor of $\infty$-categories satisfies the appropriate universal property. However, the underlying $\infty$-category of $C^{/X}$ is $(C^{th})^{/X}$ by the construction of [GHL21, Section 4.1]. By [Lur09a, Proposition 4.2.1.5], the value of the functor on underlying $\infty$-categories is equivalent to the limit of the composite:
$(C^{th})_{/X}\rightarrow(C^{th})^{/X}\xrightarrow{p^{th}}Cat_{\infty}$
where the first map is the natural map from [Lur09a, Proposition 4.2.1.5]. Thus, the construction reduces to the of the construction of Lemma 2.3 on underlying $\infty$-categories, hence the result.
@@
Suppose that $C$ is an $\infty$-bicategory. Then if $C$ has a powering $(-)^{\Delta^{1}}$ and $C^{th}$ has all (unweighted) limits, then $C$ has all (weighted) limits
@@
By the representable nature of limits, we are reduced to proving the statement for $CAT_{\infty}$. In this case, by Lemma 2.4, we want to show that weighted limits in the underlying $\infty$-category can be computed as ordinary limits. Since $Cat_{\infty}$ is cotensored over itself this follows from the formulas for weighted colimits from [GHN17] in terms of the twisted arrow category.
@@
Suppose that $C$ is an $\infty$-category. Suppose that $C^{th}$ has all finite limits and a cotensoring $(-)^{\Delta^{1}}$. Then $C$ has all finite weighted limits.
@@
The representable nature of limits allows us to reduce to the case of weighted limits on the underlying $\infty$-category of $C$.
@@
Since $\Delta_{/J}$ is finite when $J$ is, finite weighted limits in the underlying $\infty$-category exist in the presence of ordinary finite limits and cotensors with finite $\infty$-categories. However, the latter exist by the arguments of 2.3.